The MarVoSym F nt Package

Thomas Henlich (thenlich@rcs.urz.tu-dresden.de )

2nd May 2000

This document describes how to use the Martin Vogel’s Symbols font in your L^{A}TEX documents. The font is a PS font, so you need a PS printer or ghostscript. Alternatively, you can use pdflatex to produce a PDF file. The font provides the “official” Euro currency symbol, Euro symbols which go well with the Times / Helvetica / Courier fonts, warning and safety-related symbols, astronomy symbols, zodiac signs and many others.

1 Installation

In the following, replace $TEXMF with your actual texmf directory, e. g. /usr/share/texmf.

• Copy marvosym.pfb into a directory where dvips can find it, e. g. $TEXMF/fonts/

type1/public/marvosym/.

• Copy fmvr8x.tfm into a directory where TEX will find it, e. g. $TEXMF/fonts/

tfm/public/marvosym/.

• Add the line

fmvr8x␣MarVoSym␣<marvosym.pfb

to your psfonts.map file, which should reside in $TEXMF/dvips/config/, or /var/lib/texmf/dvips/config/. If you want to make the font available to pdflatex as well, repeat the procedure with the file /var/lib/texmf/dvips/config/ pdftex.map. (Note: ␣ denotes a space character.)

• Copy marvosym.sty to a TEX input directory, e. g. $TEXMF/tex/latex/misc/.

• Optionally copy marvosym.afm into an afm-directory, e. g. $TEXMF/fonts/afm/ public/marvosym/ (TEX doesn’t need this file.)

2 Usage

In your document’s preamble, type \usepackage{marvosym}. To insert a symbol, use a macro as described in the next section. The symbol will appear in the currently selected font size. To get a bigger symbol, use a size-changing command like

{\small\Smiley}{\Large\Smiley}{\LARGE\Smiley}:

3 Available symbols

3.1 Communication

\Pickup     \Letter     \Mobilefone     \Telefon     \fax     \FAX     \Faxmachine     \Email     \Lightning     \Emailct

3.2 Engineering

   \Beam         \Bearing      \Loosebearing      \Fixedbearing       \Lefttorque         \Righttorque          \Lineload          \Force

\Octosteel     \Hexasteel      \Squaresteel        \Rectsteel         \Circsteel          \Flatsteel          \Squarepipe       \Rectpipe

  \Lsteel        \TTsteel        \Circpipe           \Tsteel        \RoundedTsteel      \RoundedTTsteel      \RoundedLsteel

3.3 Information

 \Industry       \Coffeecup      \Rightscissors        \Kutline       \Leftscissors       \Football       \Bicycle

   \Info         \Clocklogo         \Cutright          \Cutline          \Cutleft        \Wheelchair     \Gentsroom     \Ladiesroom

\Checkedbox     \Crossedbox      \Pointinghand      \Writinghand

3.4 Laundry

\WashCotton     \WashSynthetics       \WashWool         \Handwash          \Dontwash            \Tumbler          \NoTumbler      \NoChemicalCleaning

  \Bleech           \NoBleech         \CleaningA        \CleaningP        \CleaningPP          \CleaningF         \CleaningFF

 \IroningI         \IroningII        \IroningIII       \NoIroning       \AtNinetyFive      \ShortNinetyFive        \AtSixty           \ShortSixty

\ShortFifty         \AtForty         \ShortForty     \SpecialForty      \ShortThirty

3.5 Currency

\EUR    \EURdig     \EURhv     \EURcr     \EURtm     \Ecommerce     \Shilling     \Denarius     \Pfund     \EyesDollar

                                                                                 \Deleatur

• Hey, \Ecommerce is not really a currency symbol, you might say. But it has something to do with money, so there you go…

• The \Denarius symbol is also known as the correction sign “Deleatur”.

• \EUR is the normal (natural) width Euro symbol. \EURdig has “special” metrics, so it has the same width as the digits (of this font). (To line up properly in tables etc.)

3.6 Safety

\Stopsign     \CEsign     \Estatically      \Explosionsafe      \Laserbeam     \Biohazard     \Radioactivity      \BSEfree

3.7 Navigation

\RewindToIndex      \RewindToStart      \Rewind     \Forward     \ForwardToEnd      \ForwardToIndex      \MoveUp     \MoveDown     \ToTop     \ToBottom

3.8 Computers

\ComputerMouse      \SerialInterface      \Keyboard     \SerialPort     \ParallelPort      \Printer

3.9 Maths

   \MVZero             \MVOne             \MVTwo         \MVThree         \MVFour       \MVFive     \MVSix     \MVSeven     \MVEight     \MVNine

\Corresponds      \Vectorarrowhigh      \Anglesign     \Vectorarrow      \Squaredot

3.10 Biology

\Neutral     \Male     \Hermaphrodite      \Female     \MALE     \HERMAPHRODITE      \FEMALE     \MaleMale     \FemaleFemale      \FemaleMale

3.11 Astronomy

\Sun    \Moon     \Mercury     \Venus     \Mars     \Jupiter     \Saturn     \Uranus     \Neptune     \Pluto     \Earth

3.12 Astrology

\Aries     \Taurus     \Gemini     \Cancer     \Leo    \Virgo     \Libra     \Scorpio     \Sagittarius      \Capricorn     \Aquarius     \Pisces

3.13 Others

\Yinyang     \Rightarrow       \MVAt      \FHBOlogo     \FHBOLOGO     \FullFHBO       \Mundus          \Cross        \Celtcross     \Ankh

 \Heart       \CircledA      \Bouquet      \Frowny       \Smiley         \Bat       \Womanface     \MartinVogel

4 Authors

The font was designed by Martin Vogel (martin.vogel@fh-bochum.de). See http://www.marvosym.de/ .

The macros and this documentation were written by Thomas Henlich (thenlich@rcs.urz.tu-dresden.de , who also converted the font to a Type 1 font. The latter involved running ttf2pt1 and doing some manual fixes afterwards.

5 History

2000-04-21 Updated the font and documentation. Changed /FontName to MarVoSym. Many

          new glyphs. Removed: \Kross, \Snowflake, \Circles, \Womanside,
          \Manside, \Womanfront, \Manfront.

1998-07-20 Changed (*) to /* in /FontName. Thanks to Denis B. Roegel for telling me about this.

1998-06-21 Conversion to type 1 font now done with ttf2pt1 program. Font works now with dvips 5.78 and partial font downloading. Thanks to Uwe W. Gehring and Armin Geisse for cooperation. Added \Ankh macro. Renamed some macros.

1998-06-10 First version.

6 Software

ttf2pt1, the TrueType to PS type 1 font converter, is free software. See http://ttf2pt1.sourceforge.net/ for more information.